A shopkeeper buys a number of packets of biscuits for Rs.80.If had bought 4 more packets for the same amount,each packet would have lost Rs.1 less.How many packets he did buy?
Answer by sachi(548)

let he buys x packets for Rs 80
so cost of each packet =Rs 80/x
now he bought 4 more packets for the same amount,
so each packet cost=Rs 80/(x+4)
as per question 80/x-80/(x+4)=1
or 80[x+4-x]/[x(x+4)]=1
or 80*4=x(x+4)=x^2+4x
or x^2+4x-320=0
or x^2+20x-16x-320=0
or x(x+20)-16(x+20)=0
or (x+20)(x-16)=0
or x=-20 or 16
taking the positive value he bought 16 packets
ans